# Personality Pairing Improves Human-AI Collaboration

### v1

@@ Results
We also evaluated the real world performance of the ads produced by our human-AI teams in a large-scale field experiment on X. Our field study evaluated traditional marketing metrics, including click-through rates (CTR) and cost-per-click (CPC), as well as downstream ad behaviors, including view-through rates (VTR) and view-through duration (VTD), using a unique DocSend link for each ad (Figure 5). VTR and VTD were measured to assess whether participants created click-bait ads irrelevant to the research report, which could elevate CTR but reduce VTR and VTD. In line with previous findings (ju2025collaborating), our mixed effects regression analysis revealed that higher quality ad text reduced CPC while higher quality images improved CTRs.
Figure 5: Regression coefficients for human-AI personality interaction terms and ad quality ratings on click-through rates (CTR; basis points) and cost-per-click (CPC; \$) and view-through rates (VTR; %) (p<<0.05). Higher values indicate better performance for CTR and VTR, whereas lower values indicate better performance for CPC. Error bars show 95% confidence intervals. See SI Section K for the full regression tables.
Personality pairings also significantly influenced ad performance. For conscientious humans, neurotic AI lowered CTRs, while for neurotic humans, neurotic AI increased CTR performance. Conscientious humans with neurotic AI and extroverted humans with agreeable AI both experienced significantly higher CPC. Open humans experienced lower VTRs when paired with conscientious AI, while conscientious humans produced ads with higher VTRs when paired with open AI. We found no statistically significant effects for VTD. Together, these results indicate that human-AI personality pairings influence ad performance both by shaping quality metrics and through effects on ad performance not captured by quality measures.
Productivity
In addition to evaluating the quality of the ads produced, we examined the effects of personality pairings on the productivity of human-AI teams. We defined productivity as the number of ad submissions completed by participants within a 40-minute task period. Since participants were given explicit monetary incentives to produce a greater number of ads and to produce higher quality ads, this definition offers a quantitative measure of output, and thus productivity, in the collaborative process of creating display ads.
Our analysis revealed several significant effects of personality pairing on productivity (Figure 6, see SI Section H). For extroverted humans, agreeable AI agents increased productivity, and for agreeable humans, open AI increased productivity while neurotic AI decreased productivity. These findings indicate that specific personality pairings influence productivity in human-AI collaboration. In particular, neuroticism mediates how agreeable, or sycophantic, AI impacts productivity. There were no significant average treatment effects for AI personality alone (see SI Section J).
Figure 6: Regression coefficients for human-AI personality interaction terms for productivity (i.e., the number of submissions; p<<0.05) for three human personality covariates, extroversion, agreeableness, and neuroticism. Error bars show 95% confidence intervals. See SI Section H for the full regression table.
Interestingly, some personality pairings introduced a trade-off between quality and productivity. Pairing agreeable humans with open AI increased productivity but decreased image quality. Pairing agreeable humans with neurotic AI had the opposite effect—they produced fewer ads but with higher text quality. These results suggest a trade-off in human-AI collaboration whereby some personality pairings improve the productivity of human-AI teams in terms of the volume of output produced per unit time, while other pairings improve ad quality while degrading productivity.
@@ Discussion
We provide the first experimental evidence that personality alignment improves human-AI collaboration and team performance. Our work highlights the need to treat AI agents not just as tools, but as teammates whose traits can be intentionally engineered for collaborative success. Recently, others have tailored human approaches to working with AI, for example by shaping how salespeople interact with AI under different decision-making authority, training, and incentives conditions, to improve sales performance (krakowski2025human). To our knowledge, ours is the first study to measure the productivity, performance and teamwork effects of AI personalization in a context in which AI personalities are directly and randomly manipulated. As AI agents become more prevalent in collaborative settings, designing them with personality traits that complement their human counterparts could enable optimization of productivity and performance in real agentic workflows and unlock more effective and satisfying collaborations.
Our results extend the person-team fit literature to human-AI teams and reveal key patterns in human-AI collaboration dynamics (barrick1998gma, bell2007meta). For example, conscientious human-AI pairs demonstrated higher effort and cohesion, suggesting that shared conscientiousness creates a structured collaborative environment that promotes more effective interaction. In addition, agreeable AI, which is in essence sycophantic AI (sharma2025sycophancy), increased productivity for extroverted and neurotic people, but reduced all quality ratings when paired with conscientious people. This suggests that sycophantic or agreeable AI enables certain humans to produce more output per unit time, but erodes the quality of that output, perhaps because it pushes back less against human work and therefore fails to improve the work while letting a greater amount of lower quality work through to production. While agreeable AI can lead to over-reliance on AI in psychosocial contexts (fang2025psychosocial), our results suggest they can also be tuned to shape performance and productivity when tailored for individual human characteristics. This suggests that sycophantic AI could be managed toward productive outcomes when applied in the right contexts for the right people and avoided in the wrong contexts and for the wrong people. These findings highlight how different personality alignments activate distinct collaborative mechanisms, with some improving teamwork and others increasing productivity. The variation in productive outcomes associated with personality matching further demonstrates that effective human-AI collaboration requires strategic consideration of trait compatibility rather than a one-size-fits-all approach.
To make the practical implications of these results concrete, consider several vignettes that suggest particular pairing choices for conscientious, agreeable, extroverted and neurotic human workers. On one hand, conscientious people benefit from conscientious and open AI. On the other hand, agreeable people are more productive but produce lower quality work when working with open AI and produce higher quality work at a slower rate when working with neurotic AI. Extroverts benefit from extroverted AI, but produce lower quality work when collaborating with conscientious AI. Neurotic people perform better when paired with extroverted AI, but worse when working with conscientious AI. These vignettes highlight concrete design choices for matching AI personalities to workers to improve human-AI collaboration.
Although our study is the first randomized control trial of AI personalization in a real-world work task and provides the first empirical evidence that human-AI personality alignment can improve collaboration, productivity and performance in human-AI teams, it is not without limitations. First, the exact personality pairing effects discovered in our marketing context may not extend to other collaborative tasks, such as coding or analytics. Indeed, our analysis of the productivity and performance effects of human-AI personality alignment in a multi-modal ads creation task, which involved both visual ad images and textual ad copy, itself revealed that personality pairing effects can vary across tasks. Future research is therefore needed to measure pairing effects and their generalizability or jaggedness across tasks.
Second, our personality prompts were independently randomized and analyzed as interaction effects with human personalities. As such, we did not examine the performance effects of any optimization procedure or intentional pairing of human to AI personality traits. In this way, our results provide a rigorous, large-scale existence proof for the potential of optimized AI personalization, rather than exploring hypothesis driven, intentional or strategic personalization per se. While our extensive randomization allows us to analyze a wide range of diverse effects to efficiently generate hypotheses, this approach requires a large sample to have sufficient statistical power and becomes challenging to analyze in a theoretically driven manner. Future studies could focus on specific personality pairings to directly and efficiently answer theoretical hypotheses or evaluate optimized personalization, especially given the hypotheses generated from this study. Furthermore, while we employ a version of P2 prompting (jiang2023evaluating), new and developing methods may better steer AI toward desired personalities (chen2025persona), and these methods should also be investigated as avenues toward strategic AI personalization.
Third, our focus on the Big Five personality traits, though robust and widely used in human personality research, may overlook other personality traits potentially more relevant to human-AI dynamics. Our analyses indicate that demographics, like country of origin, moderate performance effects, suggesting broader heterogeneity in trait interactions that warrant further exploration. Future work could also investigate linguistic patterns to better understand traits that influence teamwork (wan2025personalized).
Fourth, the study’s single 40-minute session also limits understanding of how these relationships develop over time, such as over-reliance on AI (fang2025psychosocial). Overall, these considerations highlight avenues for future research, including in-depth textual and interaction analyses, to expand on this study’s contributions to human-AI collaboration.
Finally, our study says little about the psychosocial implications of AI personalization. Although our teamwork variables address self-assessed teamwork dynamics, we examine no measures of well-being or outcomes related to, for example, AI psychosis. As the field of AI personalization develops, it will be important to understand the psychological, social, and psychosocial implications of tailoring AI personalities to their human users and we encourage the development of that line work in future research.
@@
Taken together, the effects shown here underscore the critical role of personality alignment in improving human-AI collaboration. By demonstrating that specific pairings can improve performance while others hinder it, we highlight the nuanced dynamics at play in human-AI teams. The observed trade-offs between quality and productivity further illustrate the complexity of these interactions. These insights pave the way for designing AI agents that are not only technically proficient but also socially and behaviorally adaptive, ensuring they complement their human counterparts effectively. Ultimately, our work demonstrates that personality-aligned AI has the potential to transform collaborative workflows and performance in human-AI teams.
@@ Methods
This research was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects (see osf.io/jfzha and osf.io/95dhu).
@@
The study was conducted on the Pairit platform (Figure 7). On the left panel of the platform, users can read the task, select images from a carousel, generate images using Dall-E, and write copy for the display ads. In the human-AI teams, the AI has full context of all elements on the platform, including screenshots of the ad images as they evolve over time during the collaboration. In all conditions, only the participants can submit ads. Additional details on the platform are in SI Section A-B.
Figure 7: The Pairit platform. The task workspace is on the left, and the messaging interface on the right. The participant chats with an AI agent which has full context of the user interface and can modify all elements on the task panel.
To enable the AI agents to act as effective collaborators on the Pairit platform, we provided them with full context of the task and collaboration by prompting each LLM call with all information on the user interface. The prompts included the same text of the task given to participants, previous submissions, personality prompts, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, general instructions, and a screenshot of the image after each change by the participant. We used OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06). Additional details in the AI agent are in SI Section  D.
@@
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting using OpenAI’s structured output feature (wei2023chainofthought). Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message.
For the randomization of AI personalities, we prompt the model with prompts that induce high or low levels of the Big Five personality traits. Each of the five traits is independently randomized to be high ($p=0.5$) or low ($p=0.5$; Figure 1B) using P2 prompting, which generates a detailed description of individuals with the traits (jiang2023evaluating). See SI Section E for covariate balance of the randomization of each personality prompt.
@@
As soon as participants were redirected to our platform from Prolific, they were randomized into either the human-human or the human-AI condition (Figure 1A). In the human-human condition, which we do not discuss in this study, participants joined a queue until another participant was available, at which point they were paired with each other. In the human-AI condition, which is the focus of this study, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey. Before the participants enter the task, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (rammstedt2007personality). See SI Section E for summary statistics and covariate balance of the human-AI randomization and SI Section G for the survey items.
@@
Upon completion of the task, participants completed a 35-item survey assessing teamwork quality and perception of AI (hoegl2001). The original instrument included 38 questions on a 7-point Likert scale, but three communication-related items were excluded as they were irrelevant to a one-time online collaborative effort. Next, the survey included two items on prior AI experience (i.e., “I have previously used AI chatbots (e.g., ChatGPT, Bard),” and “My experiences with AI chatbots have been positive”), one item on whether they perceived their partner as AI (i.e., “I thought my collaborator was an AI during the task”), and a final item disclosing the partner’s identity (human or AI) and asking if this altered their view of the collaboration quality (i.e., “Your collaborator was an AI/a human. How has this knowledge affected your perception of the collaboration’s quality?”). 85% of participants completed the post-task survey.
@@
To incentivize the production of high-quality ads, we informed the participants of their eligibility for additional \$100 as bonus payment, determined by the volume and quality of their submissions. Participants were explicitly instructed that “the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each.
@@
To estimate ad performance, we created and used mockups of display ads in the survey (see SI Section C). For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad."
@@
We evaluated real-world ad performance on the X platform, pre-registered at osf.io/95dhu. Key metrics included click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; fraction of document viewed), and view-through duration (VTD; seconds). VTR and VTD were tracked via a unique DocSend link for each ad. Our ad campaigns assessed causal ad effects conditional on online algorithms and divergent targeting (braun2024adplatforms, braun2024abtesting). No holdout group was used, as unique DocSend links imply zero visits without ads (braun2024adplatforms).
We sampled 2,000 ads from 11,138, selecting 1–2 ads per 1,834 human-human and human-AI teams using stratified sampling. Ads were grouped by team type and divided into 10 strata based on human-predicted click likelihood. Eight ads were excluded for potential policy violations (e.g., violence, sexual, or drug content). To avoid audience overlap, 400 campaigns each ran five unique ads as split tests. Each campaign targeted 133 random ZIP codes with population between 10,000 and 100,000, as per the 2020 American Community Survey. One-way ANOVA yielded non-significant results across ZIP codes: $F(399,53199)=0.954$, $p=0.734$ for population; $F(399,53199)=0.973$, $p=0.636$ for income. Ads were run from January 21 to February 9, 2025, with 50 campaigns active for two days each due to platform constraints. Campaign random effects controlled for temporal confounders. In our regression analyses, advertising spend was controlled for to account for auto-bidding.
@@
To evaluate the effect of personality trait interactions on human-AI team outcomes, we use the following regression model:
$Y_{J}=\sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}+\epsilon_{i}\text{,}$ (1)
where $Y_{J}$ is the outcome for observation $i$, BF are the Big Five traits, $h\in\text{BF}$ is a human trait, $a\in\text{BF}$ is a prompted AI trait, $\beta_{h}$ and $\beta_{a}$ are coefficients for the main effects of human and AI traits respectively, $\beta_{ha}$ captures the interaction between human trait $h$ and AI trait $a$, $T_{h,i}$ is the normalized (0 to 1) score on trait $h$ for the human participant associated with observation $i$, $T_{a,i}$ is an indicator equal to 1 if the AI partner was prompted to exhibit high levels of trait $a$ (0 for low), and $\epsilon_{i}$ is the error term. The unit of observation $i$ varies by outcome: for ad quality ratings, it is the individual rating (with standard errors clustered at the ad level); for productivity, teamwork quality measures, and message type fractions, it is the individual participant (with heteroskedasticity-robust standard errors). The outcomes of interest $Y_{i}$ include human-evaluated ad quality, productivity (i.e., the number of ads submitted), and teamwork quality measures.

### v2

@@ Results
Having established the effects of personality pairing on human and AI rated ad quality, we evaluated the real-world performance of the ads in a large-scale field experiment on X (Braun et al., 2024, Braun and Schwartz, 2024). The field experiment measured click-through rates (CTR), cost-per-click (CPC), view-through rates (VTR), and view-through duration (VTD) using unique DocSend links for each ad. The quality of the ads predicted their field performance as higher quality ad text reduced the cost per click (CPC) of the ads, while higher quality images increased the click through rate (CTR) on the ads (Figure 4A). These results trace the impact of personalization on performance, as personality pairings affect ad quality and ad quality directly affects ad performance in the field.
Figure 4: Effect of personality pairing on field performance. Mixed-effects regression coefficients with campaign random effects. Panel A shows the effect of ad quality (image, text, click likelihood) on click-through rate (CTR) and cost-per-click (CPC); symbols reflect raw $p$-values from the quality-only regression. Panel B shows human-AI personality interaction coefficients on CTR after controlling for ad quality (image, text, click, spend); subplot titles show the full human trait names and x-axis labels use O = Openness, C = Conscientiousness, E = Extraversion, A = Agreeableness, and N = Neuroticism. Panel B symbols indicate significance after BH correction within the 25-interaction CTR testing family: $^{\dagger}p<0.10$, $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$. Error bars show 95% confidence intervals. See SI Appendix, Section LABEL:si:full, for the full field regression tables and additional outcomes (CPC, VTR, VTD), and Section LABEL:si:robustness for BH corrections across all testing families.
Personality pairings also influenced ad performance beyond what ad quality alone could explain (preregistered; see Materials and Methods and SI Appendix, Section LABEL:si:prereg). Even when controlling for image quality, text quality, and click likelihood, neurotic humans paired with neurotic AI achieved significantly higher click-through rates (Figure 4B; $p=0.036$; SI Appendix, Section LABEL:si:robustness). Other pairings shifted costs-per-click and view-through rates, though these additional effects did not survive multiple hypothesis testing corrections. These results suggest that personality pairings also influence ad performance through channels not fully captured by our ad quality measures alone (SI Appendix, Section LABEL:si:full). Interestingly, the same neurotic-neurotic pairing was also the only pairing that significantly elevated delegation behavior and directive language in human-AI collaboration (SI Appendix, Section LABEL:si:interaction-content), a process feature shown in prior work to predict click ratings in human-AI teams (Ju and Aral, 2025).
@@ Discussion
Our study provides the first large scale experimental evidence that personality pairing improves human-AI collaboration and team performance and highlights the need to treat AI agents not just as tools, but as teammates whose traits can be intentionally engineered for collaborative success. Recently, others have tailored human approaches to working with AI, for example by shaping how salespeople interact with AI under different decision-making authority, training, and incentives conditions, to improve sales performance Krakowski et al. (2025). To our knowledge, ours is the first study to measure the performance effects of AI personalization in a context in which AI personalities are directly and randomly manipulated. As AI agents become more prevalent in collaborative settings, designing them with personality traits that complement their human counterparts could enable optimization of performance in real agentic workflows and unlock more effective and satisfying collaborations.
Our results extend the person-team fit literature to human-AI teams (Barrick et al., 1998, Bell, 2007, Kristof-Brown et al., 2005). The finding that pairings between different traits outperform same-trait pairings is consistent with complementary fit, in which team members fill each other’s gaps, rather than supplementary fit based on similarity (Kristof-Brown et al., 2005). The strongest effects were negative: extraverted humans produced lower quality ads when paired with conscientious AI, and conscientious humans produced lower quality ads when paired with agreeable AI. These results suggest that certain trait combinations create friction that impairs creative output, rather than a universal benefit of personality matching.
The findings point to actionable pairing principles. For ad quality, extraverted humans produced lower-quality ads when paired with conscientious AI (the strongest effect in the study), conscientious humans produced lower-quality ads when paired with agreeable AI, and neurotic humans produced lower-quality ads when paired with conscientious AI. For field click-through rates, neurotic humans achieved higher CTRs when paired with neurotic AI, an effect that persisted even after controlling for ad quality.
Moreover, the neurotic-neurotic field result aligns with how the team actually communicated. Among the 25 personality pairings, this was the only one that raised the pooled rate of directive language significantly ($p=0.002$, driven by both speakers), while no pairing affected message length (SI Appendix, Section LABEL:si:interaction-content). Our prior work (Ju and Aral, 2025) reports that directive, task-oriented communication is positively associated with click ratings in human-AI teams. The pairing that creates the highest increase in click-through rates is the pairing that produces the communication pattern independently shown to predict higher click through rates. Although we treat this convergence of results as descriptive rather than as evidence of a mediating effect, it suggests that the field effect operates through the kind of communication exchange two neurotic partners generate together and gives us greater confidence in the substantive accuracy of our results.
While ours is the largest randomized controlled trial of AI personalization to date, several limitations circumscribe our conclusions. First, the specific pairing effects we identified in a marketing context may not extend to other collaborative tasks such as coding or analytics. More work is needed to understand AI personalization across tasks, especially given the jagged frontier of AI capabilities, which could simultaneously drive and interact with personality effects.
Second, we randomized AI personality across the full design rather than preregistering a narrow set of targeted pairings. A focused preregistered replication of a small set of pairings would be an excellent next step, and in principle a stronger test of any particular pairing. At the same time, such a study requires knowing in advance which pairings are worth targeting. We therefore view our work as establishing that personality pairing can matter for human-AI collaboration and field performance, and as a foundation for future work on targeted replication and personalization optimization.
Third, our results were obtained with a specific model snapshot (gpt-4o-2024-08-06). Different architectures and alignment procedures produce different behavioral responses to personality prompts, so the specific coefficients are likely model-dependent. The general finding that personality pairing matters should still generalize to any model capable of prompt-induced behavioral variation. Steering techniques themselves generalize across both models and methods, ranging from prompting to representation-level interventions (Zou et al., 2023). Diversifying AI personas through personality prompts has been shown to eliminate the homogenization of creative outputs observed when all users interact with identically prompted AI (Wan and Kalman, 2026). These prior results suggest that personality-based personalization is a robust design lever rather than an artifact of any single model or technique.
Fourth, the Big Five personality traits may overlook other traits more relevant to human-AI dynamics. Our exploratory analyses point to further heterogeneity in demographic moderators like country of origin and experience with AI (SI Appendix, Section LABEL:si:demographics, and linguistic patterns in collaboration offer another promising direction for future work on dimensions relevant to AI personalization (Wan et al., 2025).
Fifth, our scope limits what we can say about longer time horizons, larger teams, and psychosocial effects. A single 40-minute session cannot address how pairings play out over repeated interactions, including whether prolonged use produces over-reliance on AI (Fang et al., 2025). We focus on one-to-one pairings rather than multi-person teams or networks of collaborators, where AI could vary its personality across teammates or over time. We measure self-assessed collaboration dynamics with teamwork variables, but we examine no measures of well-being or adverse psychosocial outcomes. As the field of AI personalization develops, it will be important to understand the psychological and social implications of tailoring AI personalities to their human users, and we encourage future research in this area.
@@
While acknowledging these limitations, we believe this large-scale experimental evaluation of personality pairing in human-AI collaboration is an important first step toward robust AI personalization. The effects shown here underscore the critical role of personality pairing in improving human-AI collaboration. Specific pairings improved ad quality, teamwork, and real-world ad performance, while other pairings hindered these outcomes. These insights pave the way for designing AI agents that are not only technically proficient but also socially and behaviorally adaptive, ensuring they complement their human counterparts effectively. Ultimately, our work demonstrates that strategically pairing AI personalities with human traits has the potential to transform collaborative workflows and performance in human-AI teams.
@@ Methods
This research was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects (see osf.io/jfzha and osf.io/95dhu). Informed consent was obtained from all participants prior to their participation in the study.
@@
The study was conducted on the Pairit platform (Figure 5). On the left panel of the platform, users can read the task, select images from a carousel, generate images using Dall-E 3, and write copy for the display ads. In the human-AI teams, the AI has full context of all elements on the platform, including screenshots of the ad images as they evolve over time during the collaboration. In all conditions, only the participants can submit ads. Additional details on the platform can be found in the SI.
Figure 5: The Pairit platform. The task workspace is on the left, and the messaging interface on the right. The participant chats with an AI agent which has full context of the user interface and can modify all elements on the task panel. Some text in the screenshot is intentionally blurred to anonymize identifying information about the partner think tank, and a larger version of the same screenshot is available in SI Appendix, Section LABEL:si:pairit.
To enable the AI agents to act as effective collaborators on the Pairit platform, we provided them with full context of the task and collaboration by prompting each LLM call with all information on the user interface. The prompts included the same text of the task given to participants, previous submissions, personality prompts, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, general instructions, and a screenshot of the image after each change by the participant. We used OpenAI’s multimodal gpt-4o model (specifically the gpt-4o-2024-08-06 snapshot) with a temperature of 0.2 and default nucleus sampling parameters; no top-$p$, max token, or seed constraints were imposed, so sampling was stochastic. Image generation used dall-e-3 at $1024\times 1024$ resolution. Additional details on the AI agent are in SI Appendix, Section LABEL:si:methods.
@@
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting using OpenAI’s structured output feature (Wei et al., 2022). Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message.
Personality prompts
We note that these prompts induce behavioral styles rather than latent psychological traits; we use “personality” throughout as shorthand for prompt-induced behavioral tendencies.
For the randomization of AI personalities, we prompt the model with prompts that induce high or low levels of the Big Five personality traits. Each of the five traits is independently randomized to be high ($p=0.5$) or low ($p=0.5$; Figure 1B) using P2 prompting, which generates a detailed description of individuals with the traits (Jiang et al., 2023). See SI Appendix, Section LABEL:si:randomization, for sample summary statistics, balance across the human-AI and human-human conditions, and the independent AI-trait randomization design.
We validated that the personality prompts produced measurable behavioral differences in AI messages. For each trait, we extracted a pre-specified set of linguistic features from AI-generated messages and constructed composite behavioral indices using the sum of standardized feature values with theoretically expected signs (Anderson, 2008). Examples of core behavioral markers include: (1) Openness: emoji usage, conservative language (reversed), and exploratory suggestions; (2) Conscientiousness: first-message elaboration (length and word count); (3) Extraversion: enthusiasm markers, exclamation usage, and words per message; (4) Agreeableness: permission-asking ratio, affirming language, and we/our orientation; (5) Neuroticism: cautious language, fewer alternatives offered (reversed), and less confident language (reversed).
All five composites distinguished their target trait condition (high vs. low) at $p<0.05$, with Cohen’s $d$ ranging from 0.12 to 0.17 (SI Appendix, Section LABEL:si:manipulation). Agreeableness had the strongest individual behavioral feature (permission-asking; $d=0.14$, $p=0.013$), and conscientiousness had the broadest manipulation (three individually significant features). A discriminant validity matrix confirmed that each composite preferentially differentiated its own trait condition, with the largest effect on the diagonal in four of five cases. LLMs can exhibit social desirability biases that attenuate personality manipulations (Salecha et al., 2024), which may partly explain the modest effect sizes of our behavioral composites.
@@
As soon as participants were redirected to our platform from Prolific, they were randomized into either the human-human or the human-AI condition (Figure 1A). In the human-human condition, which we do not discuss in this study, participants joined a queue until another participant was available, at which point they were paired with each other. In the human-AI condition, which is the focus of this study, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey. Before the participants enter the task, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (Rammstedt and John, 2007). See SI Appendix, Section LABEL:si:randomization, for summary statistics and covariate balance of the human-AI randomization, and SI Appendix, Section LABEL:si:survey, for the rating interface and ad mockups used in the evaluation study.
@@
Upon completion of the task, participants completed a 35-item survey assessing teamwork quality and perception of AI (Hoegl and Gemuenden, 2001). The original instrument included 38 questions on a 7-point Likert scale, but three communication-related items were excluded as they were irrelevant to a one-time online collaborative effort. Next, the survey included two items on prior AI experience (i.e., “I have previously used AI chatbots (e.g., ChatGPT, Bard),” and “My experiences with AI chatbots have been positive”), one item on whether they perceived their partner as AI (i.e., “I thought my collaborator was an AI during the task”), and a final item disclosing the partner’s identity (human or AI) and asking if this altered their view of the collaboration quality (i.e., “Your collaborator was an AI/a human. How has this knowledge affected your perception of the collaboration’s quality?”). 85% of participants completed the post-task survey.
@@
To incentivize the production of high-quality ads, we informed the participants of their eligibility for an additional \$100 bonus payment, determined by the volume and quality of their submissions. Participants were explicitly instructed that “the greater the number of ads, the greater your chances, but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each.
@@
To estimate ad performance, we created and used mockups of display ads in the survey (SI Appendix, Section LABEL:si:survey). For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." We construct the ad quality index at the individual rating level as the unweighted mean of each rater’s three 7-point evaluations of a given ad: text quality, image quality, and click likelihood.
@@
We evaluated real-world ad performance on the X platform, pre-registered at osf.io/95dhu. Key metrics included click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; fraction of document viewed), and view-through duration (VTD; seconds). VTR and VTD were tracked via a unique DocSend link for each ad. Ad platforms like X use auto-bidding algorithms that optimize delivery based on expected engagement, producing divergent targeting across ads within a campaign (Braun and Schwartz, 2024). Following Braun et al. (Braun et al., 2024), we estimate causal effects of ads conditional on the platform’s optimization algorithm. The relevant estimand is the effect of ad content on performance given the audience the platform selected for each ad, which is the economically relevant quantity for advertisers operating within algorithmic environments. Viewer composition is therefore not constant across ads within a campaign, but campaign random effects absorb systematic differences in audience pools across campaigns. Advertising spend is controlled to account for differential budget allocation from auto-bidding. ZIP-code-level randomization across campaigns prevents audience spillovers. No holdout group was used, as unique DocSend links imply zero visits without ads (Braun et al., 2024).
The field experiment used a subset of the human-AI ads. We sampled 2,000 ads from the 7,266 ads produced in the human-AI lab condition using stratified random sampling based on predicted click likelihood implied by the human ratings, with 10 click-score strata spanning the ad quality distribution. These ads were then deployed in 400 geographically non-overlapping campaigns, each with a five-ad split test. The selected ads covered 51.7% of human-AI lab participants (651 of 1,258), and each included participant contributed a mean of 1.77 ads. Eight sampled ads were excluded for potential policy violations on the X platform (e.g., violence, sexual, or drug content). Each campaign targeted 133 random ZIP codes with populations between 10,000 and 100,000, as per the 2020 American Community Survey. One-way ANOVA yielded non-significant results across ZIP codes: $F(399,53199)=0.954$, $p=0.734$ for population; $F(399,53199)=0.973$, $p=0.636$ for income. Ads were run from January 21 to February 9, 2025, with 50 campaigns active for two days each due to platform constraints. Campaign random effects controlled for temporal confounders.
Multiple testing corrections
Our experimental design yields 25 human-AI personality interaction terms per outcome (5 human traits $\times$ 5 AI traits). To account for multiple testing, we group related interaction tests into coherent outcome families and apply Benjamini-Hochberg (BH) corrections within each family. These families cover ad quality, teamwork, productivity (reported in the SI), and field performance, with a separate correction for each field outcome (CTR, CPC, VTR, VTD). We use the BH correction framework because our objective is to identify a set of statistically supported findings while controlling the expected false discovery rate among reported results, rather than to minimize the probability of a single false rejection across the entire family of tests. Given the number of related hypotheses evaluated, a correction framework akin to Bonferroni would impose substantially lower power and materially increase the risk of false negatives. Since our study is intended to characterize the pattern of supported effects rather than to certify that the probability of any false rejection is below a fixed threshold, BH is the most appropriate framework. We report both raw and corrected $p$-values. The pre-registration mapping table in SI Appendix, Section LABEL:si:prereg, indicates which analyses are treated as preregistered versus exploratory.
@@
We use outcome-appropriate regression models for human-AI team outcomes. For ad quality and teamwork (7-point Likert scales), we estimate ordinal logit models to respect the ordinal nature of the outcome. For productivity (submission count), we estimate Poisson generalized linear models with heteroskedasticity-robust (HC1) standard errors to accommodate dispersion in the count distribution. OLS results are reported in the SI as a robustness check and are substantively similar. The general specification is:
$g(E[Y_{i}])=\sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}\text{,}$ (1)
where $g(\cdot)$ is the link function (identity for OLS, log for Poisson, logit for ordinal), $Y_{i}$ is the outcome for observation $i$, BF are the Big Five traits, $h\in\text{BF}$ is a human trait, $a\in\text{BF}$ is a prompted AI trait, $\beta_{h}$ and $\beta_{a}$ are coefficients for the main effects of human and AI traits respectively, $\beta_{ha}$ captures the interaction between human trait $h$ and AI trait $a$, $T_{h,i}$ is the normalized (0 to 1) score on trait $h$ for the human participant associated with observation $i$, $T_{a,i}$ is an indicator equal to 1 if the AI partner was prompted to exhibit high levels of trait $a$ (0 for low). The unit of observation $i$ varies by outcome: for ad quality ratings, it is the individual rating; for productivity, teamwork quality measures, and message type fractions, it is the individual participant (with heteroskedasticity-robust standard errors). All regressions are estimated within the human-AI condition, so the human-human condition is not part of the estimation sample. Each interaction coefficient $\beta_{ha}$ in Figure 3 therefore captures how the association between human trait $h$ and the outcome differs when the AI partner is assigned high rather than low levels of trait $a$, conditional on the human and AI main effects and the other regressors. For example, if $\beta_{ha}=0.5$ for $_{\text{H}} ×\times Conscientiousness$$_{\text{AI}}$, then moving from the minimum to the maximum normalized human conscientiousness score increases the model’s linear predictor by $0.5$ more when paired with high-conscientiousness AI than with low-conscientiousness AI.
@@
Data availability
The data supporting the findings of this study will be made available in the paper’s data repository upon publication.